### v1

# TwinBrainVLA[Uncaptioned image]: Unleashing the Potential of Generalist VLMs for Embodied Tasks via Asymmetric Mixture-of-Transformers

Abstract
Standard Vision-Language-Action (VLA) models typically fine-tune a monolithic Vision-Language Model (VLM) backbone explicitly for robotic control. However, this approach creates a critical tension between maintaining high-level general semantic understanding and learning low-level, fine-grained sensorimotor skills, often leading to "catastrophic forgetting" of the model’s open-world capabilities. To resolve this conflict, we introduce TwinBrainVLA[Uncaptioned image], a novel architecture that coordinates a generalist VLM retaining universal semantic understanding and a specialist VLM dedicated to embodied proprioception for joint robotic control. TwinBrainVLA synergizes a frozen "Left Brain", which retains robust general visual reasoning, with a trainable "Right Brain", specialized for embodied perception, via a novel Asymmetric Mixture-of-Transformers (AsyMoT) mechanism. This design allows the Right Brain to dynamically query semantic knowledge from the frozen Left Brain and fuse it with proprioceptive states, providing rich conditioning for a Flow-Matching Action Expert to generate precise continuous controls. Extensive experiments on SimplerEnv and RoboCasa benchmarks demonstrate that TwinBrainVLA achieves superior manipulation performance compared to state-of-the-art baselines while explicitly preserving the comprehensive visual understanding capabilities of the pre-trained VLM, offering a promising direction for building general-purpose robots that simultaneously achieve high-level semantic understanding and low-level physical dexterity.
TwinBrainVLA[Uncaptioned image]: Unleashing the Potential of Generalist VLMs for Embodied Tasks via Asymmetric Mixture-of-Transformers
Bin Yu1,2,thanks: Equal contribution Shijie Lian2,4,11footnotemark: 1 Xiaopeng Lin2,5,11footnotemark: 1 Yuliang Wei1,thanks: Corresponding author Zhaolong Shen2,6 Changti Wu2,7 Yuzhuo Miao1,2 Xinming Wang2,8 Bailing Wang1 Cong Huang2,3 Kai Chen2,3,9,22footnotemark: 2 1HIT 2ZGCA 3ZGCI 4HUST 5HKUST(GZ) 6BUAA 7ECNU 8CASIA 9DeepCybo
footnotetext: Work done at Zhongguancun Academy (Beijing).
1 Introduction
The pursuit of embodied artificial intelligence has recently converged on the paradigm of Vision-Language-Action (VLA) models (Kim et al., 2024; Intelligence et al., 2025; NVIDIA et al., 2025). By grafting robotic control heads onto pre-trained Vision-Language Models (VLMs), these systems inherit the rich semantic reasoning and open-world generalization capabilities of large-scale internet data. This unified approach promises a robot that can not only recognize a "coffee mug" but also understand the instruction "place the mug on the coaster" and execute the corresponding motor commands.
However, a fundamental tension lies at the heart of current VLA architectures: the misalignment between the optimization objectives of semantic pre-training and robotic fine-tuning. Pre-trained VLMs are originally optimized for general semantic understanding and dialogue. In contrast, during VLA fine-tuning, the backbone is repurposed to provide precise, spatially-grounded conditioning for the Action Expert to enable precise robotic control. Forcing a single VLM backbone to satisfy these divergent objectives simultaneously inevitably precipitates "catastrophic forgetting"(Hancock et al., 2025b; Zhang et al., 2026; Zhou et al., 2025; Fang et al., 2025), where the model sacrifices its general linguistic capabilities to adapt to the specialized demands of robotic actuation. More critically, this catastrophic forgetting undermines the fundamental premise of the VLA paradigm, which aims to leverage the general world understanding acquired during VLM pre-training to enable generalized robotic control. However, current fine-tuning methodologies paradoxically strip the VLM of this very capability.
To resolve this conflict, we draw inspiration from the biological principle of hemispheric lateralization, where the human brain allocates distinct cognitive functions to specialized hemispheres—logic and language to the left, spatial attention and motor coordination to the right. We propose that VLA models should mimic this decoupling. By separating general semantic understanding from embodied action estimation, we cY-an achieve a system that is both cognitively capable and physically skillful.
Figure 1: Architectural comparison between Vanilla VLA and TwinBrainVLA.
In this paper, we introduce TwinBrainVLA[Uncaptioned image], a novel VLA framework that orchestrates two isomorphic VLM pathways via an asymmetric dual-stream joint training strategy. Structurally, our model consists of a "Left Brain" and a "Right Brain". The Left Brain is a frozen, pre-trained VLM that retains robust open-world knowledge and instruction following capabilities. The Right Brain, initialized with the same architecture, is fully trainable and specialized for embodied perception. These two streams are seamlessly fused through a Asymmetric Mixture-of-Transformers (AsyMoT) mechanism, which allows for joint attention over hidden states without sharing parameters. Crucially, the specialized representations from the Right Brain serve as the primary condition for our action expert utilizing flow-matching for precise action generation. This design ensures that the action expert receives spatially rich, task-aligned embeddings from the Right Brain, while the Left Brain explicitly preserves the model’s general semantic understanding capabilities.
Our contributions are threefold:
  • We propose TwinBrainVLA[Uncaptioned image], the first VLA architecture to explicitly decouple general semantic understanding and embodied perception via an asymmetric dual-stream design, resolving the training conflict inherent in single-backbone VLAs.
  • We introduce the Asymmetric Mixture-of-Transformers (AsyMoT) mechanism for the information interaction of two isomorphic VLM pathways to build the VLA backbone, and employ an asymmetric parameter freezing strategy to achieve the joint training of the dual models.
  • Extensive comparative experiments and evaluations on the SimplerEnv and RoboCasa benchmarks, demonstrates the effectiveness of the TwinBrainVLA architecture, the AsyMoT mechanism, and the proposed training strategy.
2 Related Work
Multimodal understanding. The landscape of computer vision and natural language processing has been revolutionized by Vision Language Models (VLMs) (Liu et al., 2023; Bai et al., 2025b, a). By seamlessly integrating visual encoders with Large Language Models (LLMs) through sophisticated projection layers or adapter modules (Jiang et al., 2025a), these models exhibit emergent abilities in holistic scene understanding, visual question answering (VQA), and complex math reasoning. Given the inherent limitations of general-purpose VLMs in spatial perception (Xiao et al., 2026; Dang et al., 2025; Zhang et al., 2025b; Chen et al., 2025; Yang et al., 2025c), recent studies have increasingly focused on employing post-training strategies to enhance spatial intelligence (Yang et al., 2025a; Tan et al., 2025; Zhou et al., 2026; Zhu et al., 2025; Li et al., 2025) and construct specialized embodied foundation models tailored for embodied scenarios (Team et al., 2025a; Yang et al., 2025a; Hao et al., 2025; Lin et al., 2025c).
Vision-Language-Action (VLA) Models. Building upon the zero-shot generalization and rich semantic priors of VLMs, VLA models have emerged as a scalable paradigm for embodied intelligence Pertsch et al. (2025); Zheng et al. (2025a); Wen et al. (2025); Shukor et al. (2025); Zheng et al. (2025b); Lin et al. (2025b); Cai et al. (2026). By fine-tuning pre-trained VLMs on large-scale robotic manipulation datasets, these models learn to map multimodal instructions and observations to low-level control signals (Jiang et al., 2025b; Fan et al., 2025; Zhai et al., 2025). To mitigate the degradation of general conversational skills during VLA training, recent approaches have integrated general visual-dialogue datasets (Mazzaglia et al., 2025; Yang et al., 2025b; Hancock et al., 2025a; Zhou et al., 2025) and reasoning mechanisms into the VLA framework (Zawalski et al., 2025; Yin et al., 2025; Lin et al., 2025a; Qu et al., 2025a; Ye et al., 2025; Huang et al., 2025; Guo et al., 2025; Lee et al., 2025; Huang et al., 2026).
3 Motivation: Catastrophic Forgetting in VLA Models
Standard Vision-Language-Action (VLA) models typically adopt a transfer learning paradigm. Formally, let a pre-trained Vision-Language Model (VLM) be parameterized by $\theta_{\text{pre}}$, which has been aligned on massive internet-scale image-text pairs to maximize the likelihood of generating semantic text responses $y$ given an image $v$ and instruction $x$. To adapt this backbone for robotic manipulation, the model is fine-tuned on a robotic demonstration dataset $\mathcal{D}_{\text{robot}}=\{(v,x,a)_{i}\}_{i=1}^{N}$, where $a$ represents the robot’s action (e.g., end-effector pose or joint angles). The training objective shifts to minimizing the action prediction loss $\mathcal{L}_{\text{action}}$:
$\theta^{*}=\arg\min_{\theta}\sum_{(v,x,a)\in\mathcal{D}_{\text{robot}}}-\log P(a|v,x;\theta),$ (1)
initialized with $\theta=\theta_{\text{pre}}$.
While this process effectively imparts manipulation skills—enabling the model to control servos and execute tasks—it introduces a critical side effect known as catastrophic forgetting. The distribution of $\mathcal{D}_{\text{robot}}$ differs fundamentally from the diverse visual-linguistic data seen during pre-training (Kachaev et al., 2025; Hancock et al., 2025a; Zhang et al., 2025a). As the parameters $\theta$ are updated to minimize the high-frequency spatial error in action space, the delicate semantic alignment established in $\theta_{\text{pre}}$ is often overwritten. Consequently, the resulting VLA model suffers from a severe degradation in its original capabilities: it loses the ability to follow complex open-ended instructions, fails to describe visual scenes accurately, and exhibits weakened general semantic reasoning. This phenomenon turns the VLM backbone from a "generalist" into a "specialist", sacrificing its linguistic brain to gain a robotic body.
4 Method: TwinBrainVLA[Uncaptioned image]
In this section, we present TwinBrainVLA[Uncaptioned image], a unified VLA framework designed to disentangle high-level semantic reasoning from fine-grained sensorimotor control. As illustrated in Figure 2, our architecture operates on an asymmetric dual-stream paradigm to resolve the "catastrophic forgetting" dilemma. The framework comprises three core components: (1) a frozen "Left Brain" that preserves open-world visual-linguistic knowledge; (2) a trainable "Right Brain" that specializes in embodied motor control; and (3) a Flow-Matching Action Expert that generates precise continuous actions. The two brains interact via a novel Asymmetric Mixture-of-Transformers (AsyMoT) mechanism, enabling the specialist to leverage generalist semantics without corrupting them. We detail the Dual-VLM architecture and AsyMoT fusion strategy in Sec. 4.1, describe the flow-matching policy formulation in Sec. 4.2, and finally present our joint-training objective in Sec. 4.3.
Figure 2: The framework of TwinBrainVLARefer to caption. (a) Overall Architecture. The model features an Asymmetric Mixture-of-Transformers design composed of two distinct pathways: a frozen "Left Brain" (Generalist) for semantic reasoning and a trainable "Right Brain" (Specialist) for embodied motor control. The Right Brain fuses visual, textual, and proprioceptive state inputs to provide conditioning for the Action Expert, which utilizes a Flow-Matching algorithm to denoise continuous robotic actions. (b) Asymmetric MoT Mechanism (AsyMoT). Through causal self-attention, the trainable Right Brain attends to the frozen Key-Value (KV) pairs of the Left Brain, enabling the transfer of general semantic knowledge to the robotic control policy without catastrophic forgetting.
4.1 Asymmetric Dual-VLM as Backbone
The backbone of TwinBrainVLA consists of two parallel Vision-Language Models (VLMs), denoted as the "Left Brain" ($\mathcal{M}_{L}$) and the "Right Brain" ($\mathcal{M}_{R}$). Both streams are initialized with identical pre-trained weights (e.g., Qwen3-VL series (Bai et al., 2025a)) but serve distinct roles during training.
Dual-Stream Inputs and State Encoding. The input to the system comprises a visual observation $\mathbf{I}$, a natural language instruction $\mathbf{T}$, and the robot’s proprioceptive state $\mathbf{s}\in\mathbb{R}^{d_{s}}$ (e.g., joint angles and end-effector pose). The Left Brain functions as a pure semantic generalist. It receives only the vision and language tokens:
$\mathbf{H}_{L}^{0}=[\mathcal{V}(\mathbf{I});\mathcal{T}(\mathbf{T})],$ (2)
where $\mathcal{V}$ and $\mathcal{T}$ represent the vision encoder and text tokenizer, respectively.
In contrast, the Right Brain is responsible for embodied control and must be grounded in the robot’s physical state. To achieve this, we introduce a lightweight State Encoder $\phi$, modeled as a Multi-Layer Perceptron (MLP), which projects the low-level proprioceptive state $\mathbf{s}$ into the VLM’s embedding space. These state tokens are appended to the multimodal input sequence:
$\mathbf{H}_{R}^{0}=[\mathcal{V}(\mathbf{I});\mathcal{T}(\mathbf{T});\phi(\mathbf{s})].$ (3)
This design ensures that the Right Brain explicitly attends to the robot’s current configuration, a critical requirement for closed-loop control.
Asymmetric Mixture-of-Transformers (AsyMoT). The core innovation of our backbone lies in how these two streams interact. We propose the AsyMoT mechanism to fuse general semantic knowledge into the robotic control stream. Let $\mathbf{H}_{L}^{l}$ and $\mathbf{H}_{R}^{l}$ denote the hidden states of the Left and Right Brains at layer $l$, respectively.
The Left Brain remains frozen during training. Its self-attention mechanism operates independently to preserve pre-trained general capabilities:
$\mathbf{H}_{L}^{l+1}=\text{Attn}(Q_{L}^{l},K_{L}^{l},V_{L}^{l})+\text{FFN}(\mathbf{H}_{L}^{l}),$ (4)
where $Q,K,V$ are derived from $\mathbf{H}_{L}^{l}$ using frozen projection weights.
The Right Brain is trainable and employs an Asymmetric Joint Attention mechanism. It queries not only its own context but also the semantic features from the Left Brain. Specifically, at each attention head, the Right Brain computes its Query ($Q_{R}$), while the Key ($K$) and Value ($V$) are constructed by concatenating the representations from both brains:
$K_{\text{joint}}$ $=[\text{sg}(K_{L}^{l})\ ;\ K_{R}^{l}],$ (5)
$V_{\text{joint}}$ $=[\text{sg}(V_{L}^{l})\ ;\ V_{R}^{l}],$ (6)
$\mathbf{H}_{R}^{l+1}$ $=\text{Softmax}\left(\frac{Q_{R}^{l}(K_{\text{joint}})^{T}}{\sqrt{d_{k}}}\right)V_{\text{joint}}+\text{FFN}(\mathbf{H}_{R}^{l}),$ (7)
where [;][\ ;\ ] denotes concatenation along the sequence length dimension, and $\text{sg}(\cdot)$ indicates the stop-gradient operation.
This asymmetric flow ensures a strict hierarchy: the Left Brain acts as a stable "semantic anchor" providing high-level reasoning features, while the Right Brain dynamically fuses these semantics with fine-grained proprioceptive cues to reason about spatial actions. The final hidden states of the Right Brain, $\mathbf{H}_{R}^{final}$, are then passed to the Action Expert.
4.2 Flow-Matching Action Expert
To enable high-precision, continuous control, we move beyond the discrete tokenization paradigm (Kim et al., 2024) and adopt the state-of-the-art generative policy architecture utilized in recent foundation models like Isaac-GR00T (NVIDIA et al., 2025) and $\pi_{0}$ (Intelligence et al., 2025). Specifically, our Action Expert employs the Diffusion Transformer (DiT) architecture (Peebles and Xie, 2023) and is trained via flow matching.
Architecture and Conditioning. Following the design in Isaac-GR00T, the Action Expert operates as a conditional decoder that denoises action trajectories. It takes a noisy action sequence $\mathbf{a}_{t}$ and the current timestep $t$ as input. While the policy architecture aligns with established baselines, a key distinction in our framework lies in the source of the condition. We inject the spatially-rich representations from our trainable Right Brain, $\mathbf{H}_{R}$, into the DiT via cross-attention layers. This ensures that the proven control capabilities of the DiT are guided explicitly by the embodied perception features extracted by our asymmetric backbone.
Flow-Matching Formulation. We utilize the flow-matching objective to train the policy, as demonstrated effectively in prior work (NVIDIA et al., 2025). The process models a conditional probability path flowing from a standard Gaussian prior $\mathbf{a}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ to the ground-truth action distribution $\mathbf{a}_{1}$. The vector field regresson loss is defined as:
[…]
where $v_{\psi}$ is the DiT network, $t\sim\mathcal{U}[0,1]$, and the target vector field is the straight line $\mathbf{a}_{1}-\mathbf{a}_{0}$. During inference, actions are synthesized by solving the Ordinary Differential Equation (ODE) using an Euler solver, translating the Right Brain’s understanding into smooth robotic movements.
4.3 Training Strategy
Optimization Objective. Consistent with standard VLA fine-tuning paradigms (NVIDIA et al., 2025), we train TwinBrainVLA using exclusively the robotic action objective. We do not incorporate auxiliary Next-Token Prediction (NTP) losses or mix in general vision-language datasets during this stage. The training objective is to minimize the Flow-Matching loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{FM}}(\theta_{R},\psi,\phi;\mathcal{D}_{\text{robot}}),$ (9)
where $\mathcal{D}_{\text{robot}}$ represents the robotic demonstration dataset, and $\{\theta_{R},\psi,\phi\}$ denote the trainable parameters of the Right Brain, Action Expert, and State Encoder, respectively. It is worth noting that in monolithic VLA architectures, relying solely on action loss typically leads to severe catastrophic forgetting of general semantic capabilities. However, our dual-stream design structurally immunizes the model against this degradation: the "Right Brain" is free to specialize entirely in control dynamics, while the frozen "Left Brain" implicitly safeguards the linguistic and semantic priors.
Asymmetric Update Rule. We enforce a strict parameter update policy. Let $\theta_{L}$ be the parameters of the Left Brain. During backpropagation, we set $\nabla\theta_{L}=0$. Gradients originating from the Action Expert flow backwards through the DiT ($\psi$), permeate the Right Brain ($\theta_{R}$) via the conditioning signals, and update the State Encoder ($\phi$). Crucially, at the AsyMoT fusion layer, the gradient flow is explicitly blocked (stop-gradient) at the Key-Value interface from the Left Brain. This ensures that the Left Brain acts as a stable semantic anchor, providing consistent features $K_{L},V_{L}$ without its weights being perturbed by the high-variance gradients typical of robotic control tasks.
5 Experiment
To comprehensively evaluate the efficacy of TwinBrainVLA, we conduct extensive experiments on two simulation benchmarks: SimplerEnv (Li et al., 2024c) and RoboCasa (Nasiriany et al., 2024). Our training pipeline is built upon the starVLA framework (starVLA, 2025), distributed across 16 $\times$ NVIDIA H100 GPUs. We strictly follow its default training protocols to ensure a fair comparison. More extensive simulation benchmarks and real-world robotic experiments are in progress.
5.1 Experiments on SimplerEnv
Implementation and Training Setup. To demonstrate the scalability and effectiveness of our architecture, we instantiate TwinBrainVLA with two state-of-the-art VLM backbones: Qwen2.5-VL-3B-Instruct and Qwen3-VL-4B-Instruct. Consistent with our asymmetric design, both the frozen Left Brain and the trainable Right Brain are initialized from these pre-trained checkpoints to ensure aligned feature spaces. For the training data, we utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: the Bridge-V2 and the Fractal dataset. We are actively working on scaling our training to the complete OXE dataset, and future experimental results will be reported in subsequent updates. Comprehensive implementation details are provided in Appendix A.
Evaluation. The benchmark consists of four manipulation tasks: "put spoon on towel", "put carrot on plate", "stack green block on yellow block", "put eggplant in the yellow basket". For each task, we evaluate our VLA policy using the evaluation script provided by the SimplerEnv repository (Li et al., 2024c). To mitigate randomness, we run 480 independent trials and report the mean performance (Avg@480).
Baselines. We benchmark TwinBrainVLA against a comprehensive suite of state-of-the-art VLA policies, including RT-1-X (O’Neill et al., 2024), Octo (Team et al., 2024), OpenVLA (Kim et al., 2024), RoboVLM (Li et al., 2024b), TraceVLA (Zheng et al., 2025c), SpatialVLA (Qu et al., 2025b), CogACT (Li et al., 2024a), VideoVLA (Shen et al., 2025), $\pi_{0}$ (Black et al., 2024), and $\pi_{0.5}$ (Intelligence et al., 2025). All baseline results are obtained from their original papers or official reproductions to ensure a fair and reliable comparison.
Table 1: Results of evaluating the VLA models with the WidowX robot in the SimplerEnv simulation environment. We highlight the best results in bold and the second-best results with underline.
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 OpenVLA-OFT (Kim et al., 2025) 12.5 4.2 4.2 72.5 23.4 RoboVLM (Li et al., 2024b) 50.0 37.5 0.0 83.3 42.7 TraceVLA (Zheng et al., 2025c) 12.5 16.6 16.6 65.0 27.7 SpatialVLA (Qu et al., 2025b) 20.8 20.8 25.0 70.8 34.4 CogACT (Li et al., 2024a) 71.7 50.8 15.0 67.5 51.3 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.1 0.0 16.6 62.5 27.1 $\pi_{0.5}$ (Intelligence et al., 2025) 49.3 64.7 44.7 69.7 57.1 QwenGR00T + Qwen3-VL-3B (starVLA, 2025) 87.5 50.0 29.2 54.2 55.2 QwenGR00T + Qwen3-VL-8B (starVLA, 2025) 70.8 38.9 22.2 65.3 49.3 QwenGR00T + Qwen2.5-VL-7B (starVLA, 2025) 59.2 30.8 3.3 44.2 34.4 Isaac-GR00T-N1.6-Bridge (Team et al., 2025b) 64.5 65.5 5.5 93.0 57.1 TwinBrainVLA + Qwen2.5-VL-3B-Instruct 83.3 41.7 31.5 77.1 58.4 TwinBrainVLA + Qwen3-VL-4B-Instruct 73.0 58.3 33.3 83.3 62.0
Results. The quantitative results on the SimplerEnv benchmark are presented in Table 1. Notably, despite not undergoing large-scale pre-training for robotic action prediction, TwinBrainVLA achieves state-of-the-art performance among all listed methods. Our framework demonstrates strong generalizability across different VLM families, attaining competitive success rates of 58.4% with Qwen2.5-VL-3B-Instruct and 62.0% with Qwen3-VL-4B-Instruct. The latter configuration surpasses the strongest baseline, Isaac-GR00T-N1.6 (57.1%), by a notable margin of +4.9%, validating the effectiveness of our asymmetric dual-brain architecture in bridging high-level semantic understanding and low-level robotic control.
5.2 Experiments on RoboCasa
Implementation. We train TwinBrainVLA on the Humanoid Robot Tabletop Manipulation subset from the PhysicalAI-Robotics-GR00T-X-Embodiment-Sim dataset (NVIDIA et al., 2025). All other experimental settings, including model initialization, distributed training infrastructure, and hyperparameters, remain identical to those described in Sec. 5.1.
Evaluation. We evaluate our policy on the RoboCasa GR1 Tabletop Benchmark, which comprises a diverse suite of 24 tabletop manipulation tasks. These tasks involve complex interactions with articulated objects and varying geometries, such as PnPBottleToCabinetClose (placing a bottle into a cabinet and closing the door), PnPCanToDrawerClose, and various microwave/toaster interaction scenarios. To ensure statistical reliability while maintaining consistency with our previous experiments, we report the success rate averaged over multiple independent trials (Avg@50) for each of the 24 tasks.
Table 2: Results of evaluating the VLA models with the GR1 robot in the RoboCasa Tabletop simulation environment. We highlight the best results in bold and the second-best results with underline.
[…]
PnP Bottle To Cabinet Close 51.5 46.0 26.0 62.0 74.0
PnP Can To Drawer Close 13.0 80.0 62.0 66.0 72.0
PnP Cup To Drawer Close 8.5 54.0 42.0 46.0 52.0
PnP Milk To Microwave Close 14.0 48.0 50.0 52.0 60.0
PnP Potato To Microwave Close 41.5 28.0 42.0 56.0 36.0
PnP Wine To Cabinet Close 16.5 46.0 32.0 58.0 46.0
PnP Novel From Cuttingboard To Basket 58.0 48.0 40.0 50.0 62.0
PnP Novel From Cuttingboard To Cardboardbox 46.5 40.0 46.0 44.0 46.0
PnP Novel From Cuttingboard To Pan 68.5 68.0 60.0 64.0 70.0
PnP Novel From Cuttingboard To Pot 65.0 52.0 40.0 60.0 66.0
PnP Novel From Cuttingboard To Tieredbasket 46.5 56.0 44.0 52.0 52.0
PnP Novel From Placemat To Basket 58.5 42.0 44.0 56.0 30.0
PnP Novel From Placemat To Bowl 57.5 44.0 52.0 59.0 54.0
PnP Novel From Placemat To Plate 63.0 48.0 50.0 62.0 64.0
PnP Novel From Placemat To Tieredshelf 28.5 18.0 28.0 28.0 38.0
PnP Novel From Plate To Bowl 57.0 60.0 52.0 62.0 60.0
PnP Novel From Plate To Cardboardbox 43.5 50.0 40.0 52.0 58.0
PnP Novel From Plate To Pan 51.0 54.0 36.0 56.0 56.0
PnP Novel From Plate To Plate 78.7 70.0 48.0 68.0 66.0
PnP Novel From Tray To Cardboardbox 51.5 38.0 34.0 40.0 46.0
PnP Novel From Tray To Plate 71.0 56.0 64.0 56.0 72.0
PnP Novel From Tray To Pot 64.5 50.0 44.0 60.0 56.0
PnP Novel From Tray To Tieredbasket 57.0 36.0 50.0 42.0 46.0
PnP Novel From Tray To Tieredshelf 31.5 16.0 28.0 34.0 28.0
[…]
Baselines. We compare TwinBrainVLA against several competitive baselines trained on the same dataset. Specifically, we include Isaac-GR00T-N1.6 (Team et al., 2025b), a large-scale VLA model that demonstrates strong performance on humanoid manipulation tasks. Additionally, to directly assess the advantages of our dual-brain architecture, we benchmark against two representative models from the starVLA framework (starVLA, 2025): QwenGR00T and QwenPI, both instantiated with Qwen3-VL-4B-Instruct backbones. These baselines follow similar training protocols and utilize the same dataset, enabling a fair comparison and highlighting the specific contributions of our asymmetric dual-VLM design.
Results. The quantitative results on the RoboCasa GR1 Tabletop Benchmark are presented in Table 2. TwinBrainVLA achieves the best performance across all 24 manipulation tasks, with our Qwen3-VL-4B-Instruct variant attaining an average success rate of 54.6%, followed closely by the Qwen2.5-VL-3B-Instruct variant at 53.5%. These results substantially outperform all baseline methods: our best model surpasses Isaac-GR00T-N1.6 (47.6%) by +7.0%, QwenGR00T (47.8%) by +6.8%, and QwenPI (43.9%) by +10.7%. These results reinforce our hypothesis that decoupling semantic understanding from embodied perception enables more effective learning of fine-grained manipulation skills in complex tabletop scenarios.
6 Conclusion
In this paper, we introduced TwinBrainVLA[Uncaptioned image], a novel VLA framework that addresses the fundamental tension between semantic generalization and embodied specialization in robotic learning. By drawing inspiration from hemispheric lateralization, we proposed an asymmetric dual-stream architecture that structurally decouples the visual cognitive process. Our AsyMoT mechanism effectively synergizes a frozen "Left Brain" for robust open-world reasoning with a trainable "Right Brain" for precise sensorimotor control, enabling the system to master complex manipulation skills without "catastrophic forgetting" of its general visual capabilities.
Limitation and Future Work
While TwinBrainVLA provides a powerful framework for decoupling semantic and embodied processing, several limitations remain that outline our future research directions.
A more decoupled model architecture. First, the current implementation of the Asymmetric Mixture-of-Transformers (AsyMoT) mechanism requires the Left and Right Brains to share an identical model architecture. This structural constraint ensures compatible hidden state dimensions for seamless joint attention but limits the flexibility to pair heterogeneous models—for instance, synergizing a massive, reasoning-heavy VLM with a lightweight, high-frequency control model. We are actively exploring generalized fusion mechanisms, such as learnable projection layers or cross-attention adapters, to support backbones of varying sizes and architectures.
More diverse model checkpoints. Second, both streams in our current setup are initialized from standard general-purpose VLM checkpoints (e.g., Qwen-VL). However, the field has arguably seen a surge in specialized Embodied VLMs trained specifically for robotics. Integrating these specialized models into our dual-brain paradigm—potentially initializing the "Right Brain" with an embodied checkpoint while retaining a generalist "Left Brain"—presents a promising avenue. We are currently investigating strategies to effectively adapt our framework to leverage these domain-specific post-trained weights.
Training on larger-scale data. Third, regarding data scale, our current training utilizes subsets of the Open X-Embodiment dataset. We acknowledge that the capacity of our dual-stream architecture may not be fully realized with limited data. We are in the process of scaling our pre-training to the complete dataset to further enhance the model’s robustness and downstream performance.
More benchmarks and real-robot experiment. Finally, future work will also focus on extending our evaluation to a broader range of benchmarks, including real-robot scenarios, to comprehensively assess the versatility of TwinBrainVLA.
Appendix A Training Hyperparameters
We instantiate the VLM backbones by initializing the weights from Qwen2.5-VL-3B-Instruct and Qwen3-VL-4B-Instruct. The model is fine-tuned for 40k steps on a cluster of 16 GPUs (batch size 16 per device). We employ the AdamW (Loshchilov and Hutter, 2019) optimizer initialized with a learning rate of 1e-5 and a cosine annealing schedule. System-level optimizations include DeepSpeed ZeRO-2 (Loshchilov and Hutter, 2019), gradient clipping at a norm of 1.0, and no gradient accumulation. Our training pipeline is built upon the starVLA (starVLA, 2025) framework.
Appendix B VLM Prompt Template
This section presents the prompt templates used for VLM input during the VLA training process.
System Prompt Template for Left Brain You are a helpful robot brain that can understand images and texts.
System Prompt Template for Right Brain You are a helpful robot brain that can understand images, texts, and robot states.
You will be provided with observation, an instruction, and the robot state. Take action to execute the instruction.
User Prompt Template for Left Brain <image>
Instruction: {instruction}
User Prompt Template for Right Brain <image>
Instruction: {instruction}
Robot Type: {robot_type}
Predict the next action for the robot based on the observation, instruction and robot type.